OpenStudy (anonymous):

A fireworks display is launched from a platform 10 feet above ground with an initial upward velocity of 70 feet per second. The height of the fireworks above ground after t seconds is given by the equation h = –16t2 + 70t + 10, where h is the height of the fireworks in feet and t is the time in seconds after they are launched. What is the maximum height of the fireworks display, to the nearest foot? A. 87 feet B. 174 feet C. 70 feet D. 10 feet

OpenStudy (cwrw238):

r o familiar with calculus?

OpenStudy (anonymous):

No, it's in an Algebra course.

OpenStudy (cwrw238):

oh ok i was just checking because you could use calculus to do this.

OpenStudy (anonymous):

Oh, I have no idea how to solve it. Do I substitute?

OpenStudy (cwrw238):

you can solve it by finding the maximum value of h you need to convert the function to vertex form

OpenStudy (cwrw238):

h = -16t^2 +70t + 10 = -16[ (t^2 -70t/16 - 10/16] = -16[t -35/16)^2 - 35^2/16^2 - 10/16] = -16[t - 35/16)^2 - 5.41] vertex = -16* - 5.41= 87 feet to nearest foot = highest point